P(5,2) , Q (7,4)


1
Expert's answer
2021-05-07T09:08:01-0400

PQ = "\sqrt{(5-7)^2+(4-2)^2} = \sqrt{2^2+2^2} = \sqrt{8} = 2\sqrt{2}"

Answer: "PQ = 2\sqrt{2}" .